(√ 3+√ 2)^x + (√ 3-√ 2)^x = 10 what is da value of x please explain..?

 
 
 
 

Grade:12th pass

2 Answers

Vikas TU
9 years ago
Let(√ 3+√ 2)^x = t
Rationalising--->
it will become,
1/(√ 3-√ 2)^x = t
or √ 3-√ 2)^x = 1/t
 
Now quadratic ------>
t + 1/t = 10
u can solve now,
 
plz approve!
 
 
sajid
8 years ago
 
Let(√ 3+√ 2)^x = t
Rationalising--->
1/(√ 3-√ 2)^x = t
or √ 3-√ 2)^x = 1/t
Now quadratic ------>
t + 1/t = 10
t*2+1-10t=0
solve it further